A company has determined its corporate mission is to pollute the world. An employee, hoping to impress the boss, tosses a cylindrical can full of toxic goo in the River. The can has a radius of 5 cm, is 20 cm high, and just barely floats (all but the top surface of the can is submerged). Determine the mass of the can of toxic goo.
Answer #1

The can is only just buoyant enough to float, so its average density is 1g/cc.

Its volume is = pie * r^2 * h

v = 3.14 * 5^2 * 20 = 1,571cc.

It is displacing this volume of the water, which is also 1g/cc. density.

Its mass is (1 * 1,571) = 1,571g., = 1.571kg

so the mass of the can of toxic goo is 1.571 kg